# DrawTalking: Building Interactive Worlds by Sketching and Speaking

### v3

@@ 4.1. Concept
In DrawTalking, speaking serves a dual-purpose: the user can explain concepts and tell stories, and at the same time refer to and label the objects in their sketched world with semantics (e.g. nouns, adjectives, adverbs) determining their names and behaviors.
@@ 4.2. User Interface Components
The interface (Figure 2) exposes a transcript view (Figure 2) and semantics diagram (Figure 4) to make system understanding of input transparent, quickly editable and accessible, and error-robust, as per our interface design goals (section subsection 3.3).
@@
Complementary to the sketching and language workflows, the find panel along with the transcript view enables the user to find and teleport to objects by semantic search, which helps avoid losing track of objects. Additionally, it enables a means to view, toggle, or delete the current commands and rules acting on the world simulation (See Figure 6).
@@ 4.3. Sketching & Labeling
Figure 4. Semantics Diagram: An example workflow of error-correction (changing the target ”building” from red house to yellow skyscraper). (a) spoken user input in the transcript panel. (b) The semantics diagram generated from the user’s input to visualize system’s interpretation. (c-d) Pen+touch interaction between diagram nouns and objects in the scene (including unlabeled ones) allows for re-linking to modify or correct the machine’s selections. (e) Confirm language command. If a verb is unknown for a command, the user can pick from an auto-presented list of similar verbs, or otherwise cancel. In our implementation, if a sentence produces too long a diagram to fit in the window, the user can zoom-out independently from the canvas.
@@
4.4. Language Commands & Functionality
DrawTalking interprets the structure of language input into commands built from primitives. These primitives are intended to be a small sample of possible functionality that demonstrate our working concept. (They are partly inspired by existing software and design spaces (Shi et al., 2021; Dreams, 2020; Resnick et al., 2009).) (See the Info Sheet supplemental material.) We emphasize that our contribution is the interactive way of controlling elements. Our specific application and the fidelity and comprehensiveness of the supported behaviors, are implementation details independent of the DrawTalking concept.
Verbs are actions performable by sketches and the system, either built-in or user-defined in terms of other verbs by composition. Examples of implemented verb primitives include:
  • animations (e.g. move, follow, rotate, jump, flee),
  • state changes (e.g. create, transform),
  • events (e.g. collides with, press),
  • inequalities (e.g. equal, exceed).
@@ 4.4. Language Commands & Functionality
(like on, to, under) cause verbs to exhibit different behavior, e.g. ”The dog jumps on/under the bed” impacts the dog’s final position relative to the target. All verbs can use any such prepositions as input. Prepositions also might describe the spatial relationship between objects in a command (Figure 7).
@@
Nouns, pronouns, and deixis refer to object labels and are used to pick specific objects or specify types of objects. Using deixis while selecting an object will select the object immediately; pronouns can refer-back to objects, enabling more natural sentences as commands.
For choosing specific sketches in a command:
  • choose 1 or more specific object instances: ”The” $+$ noun
  • choose all of a label: ”all” $+$ nouns
  • select a specific number of objects: ¡number¿ $+$ noun(s)
  • select a random object: ”a” $+$ noun
@@
Adjectives and adverbs are usable as labels that define properties on objects. Verbs receive adjectives as arguments that are evaluated continuously to modulate values and effects (e.g. magnitude, speed, size, distance, height). For example, ”fast”, ”slow,” ”excited” impact jump height and/or movement speed. Adjectives also disambiguate like-noun labeled-objects (e.g. ”first house” vs. ”second house”). Adjectives can be removed programmatically via negation, e.g. The thing is not fast.
Adverbs heighten adjective effects multiplicatively – e.g. ”very” and ”slightly” – and are chainable – e.g. ”very, very.” Special adjectives offer system-control: e.g. ”static” fixes sketches to the screen like a GUI element, useful for buttons, d-pads, score displays, etc.; ”visible”/”invisible” toggle sketch visibility by command or automatically when saying ””This is a ¡…¿”.
@@ 4.5. Examples
Figure 8. Windmill Simulation. Flexible process for constructing an interactive windmill built from user-defined rules, sketches, and triggers. The user can do this in any order and can try results at each step. This works on any sketch labeled ”blade.”
@@
[A mock-up showing the construction of a windmill whose blades rotate as wind sketches spawn-in and move past the blades, all scripted-in using speech commands for defining rules, along with hand-drawn sketches]A mock-up showing the construction of a windmill whose blades rotate as wind sketches spawn-in and move past the blades, all scripted-in using speech commands for defining rules, along with hand-drawn sketches; Panel 1 shows construction of the windmill with the base labeled ”windmill,” and the blades labeled ”blades.” The blades are attached to the windmill via the command (captioned below) ”I attach the blades to the windmill.” Panel 2 shows the user saving a hand-drawn wind spiral sketch so the system knows what to spawn when ”wind” is specified for a ”create” command. Panel 3 shows definition of rules for collision between wind and blades, captioned: ”When wind collides with blades, blades rotate.” and ”After wind collides with blades, blades stop rotating.” Panel 4 shows that extra ”switch” and left-side wall sketches have been drawn. Pressing the switch causes wind to spawn from the left wall and move to the right. Several clones of the previously saved wind sketch are shown moving to the right and colliding with the windmill blades, causing them to spin clockwise. Captioned are the commands to define all of this: ”When I press the switch I create wind at the wall.” ”After wind appears, wind moves right.”
Figure 9. Pong into Breakout. Left: per-player paddles and points; Middle, Right: scene reoriented, second player’s objects removed, destructible bricks logic defined and objects added in-bulk into the desired region. (Note: not all commands shown.)
@@
[Two screenshots of (left) a game of Pong created in DrawTalking and (right) the same game’s sketches repurposed and extended into the similar game of Break-out]Two screenshots of (left) a game of Pong created in DrawTalking and (right) the same game’s sketches repurposed and extended into the similar game of Break-out; Left: Pong contains 4 wall sketches and score counters for each player atop the walls. There is a start game switch in the bottom right that starts the game when pressed. Two paddles within the walls are being manipulated via touch controls. A ball sketch is shown bouncing towards the right paddle, which is about to reflect it left. The current score is 0-0. Right: the scene has been rotated vertically, and the second score counter has been deleted. The user has created several blocks filling the top part of the inner-walled area of the game area they’ve created. Some of the blocks have been destroyed by the user’s bouncing ball. Currently, the ball has bounced off the user’s paddle and is heading towards another breakabout brick. The score counter reads 25 for the total number of points earned from breaking the blocks. All of these mechanics have been created from user-specified rules in speech commands.
@@
Figure 10. Day/Night Cycle with Transforming Villagers. A sun/moon transforms into the other in-time with collision events. When the moon appears, the villagers transform into ghosts, which chase the boy. This shows event-based behavior changes.
@@
Figure 11. ”Dog Plays Fetch” as a Prototype Game Mechanic (demonstrating adjectives for dynamic logic) The dog interaction is re-imagined as an interactive game mechanic.
@@ 4.5.5. Day/Night Cycle: Transforming Objects and Periodic Collision Logic
4.6. Implementation
4.6.1. Hardware and Software
We implemented DrawTalking natively on the iPad Pro M2, with the main code written in C++ and C, bridged with Objective C++ and Swift to access platform-specific APIs. Speech recognition is on-device. A local NodeJS+Python server-side for NLP runs on a MacBook Pro. Text is continuously sent to the server where 1) a dependency-tree is created, 2) sent to the client and compiled into a key-value format for semantic roles (e.g. AGENT, OBJECT), 3) either interpreted in real-time by our custom engine to drive interactive simulations, or used to find objects by deixis. In final stages, the system looks-up and maps objects to roles in the command structure and executes it for animated effects. (See Figure 12 for a diagram of the the concrete client-server implementation and data-flow.)
4.6.2. Processing Steps
The language processing component comprises roughly 4 steps that translate a natural language data structure into a command for the application (Figure 13). For supplementary examples, see Appendix D.

### v4

@@ 4.1. Concept
In DrawTalking, speaking serves a dual-purpose, enabling simultaneous narration and interaction with the interface for storytelling and world-building: the user can explain concepts and tell stories, and at the same time draw, refer to, and label the objects in their sketched world with semantics (e.g. nouns, adjectives, adverbs) determining objects’ names and behaviors.
@@ 4.2. User Interface Components
The interface (Figure 2) exposes a transcript view (Figure 2) and semantics diagram (Figure 4) to make system understanding of input transparent, quickly editable and accessible, and error-robust, as per our interface design goals (subsection 3.3).
@@
Complementary to the sketching and language workflows, the find panel along with the transcript view enables the user to find and teleport to objects by semantic search, which helps avoid losing track of objects. (This use of speech is inspired by real-world interaction, in which we use language to talk about real, distant, invisible, imaginary, or conceptual objects.) Additionally, it enables a means to view, toggle, or delete the current commands and rules acting on the world simulation (See Figure 6).
@@ 4.3. Sketching and Labeling
Figure 4. Semantics Diagram: An example workflow of error-correction (changing the target ”building” from red house to yellow skyscraper). (a) spoken user input in the transcript panel. (b) The semantics diagram generated from the user’s input to visualize system’s interpretation. (c-d) Pen+touch interaction between diagram nouns and objects in the scene (including unlabeled ones) allows for re-linking to modify or correct the machine’s selections. (e) Confirm language command. If a verb is unknown for a command, the user can pick from an auto-presented list of similar verbs, or otherwise cancel (Figure 27). In our implementation, if a sentence produces too long a diagram to fit in-view, the user can zoom-out independently from the canvas.
[…]
[The workflow is shown by example for editing the linking between nouns in a command and sketch objects in the scene.]The workflow is shown by example for editing the linking between nouns in a command and sketch objects in the scene. a) Subtitled ”(Input), Tap to stage diagram.” Shows the spoken user input in the transcript view: ”the ape jumps on the building.” The user taps the ”language action” button with their finger to stage the command. b) is subtitled ”Optional Edit by unlinking.” It shows the semantics diagram representation of the input in a) as understood by the machine. The diagram shows the blocks ape, jump, on, building. Beneath ”ape” is a miniature proxy for an ape sketch. Beneath ”building” is a red house proxy. A finger is shown tapping the house, generating an arrow pointing from the proxy on the semantics diagram into the sketch canvas where the actual, original house drawing labeled ”building,” ”house” is located. A description reads ”Tap to locate original sketch in the canvas with a helpful arrow.” c) Subtitled ”Optional Edit by Linking.” A finger is shown tapping the red building on the diagram and simultaneously tapping the word ”building” with the pen to unlink the building from the diagram. A description reads ”Pen+touch sketch-to-word to remove sketch from command.” d) Pen+touch between another yellow sketch in the canvas labeled ”skyscraper,” ”building” and the word ”building” is shown, indicating that now a different sketch is being linked into the command to replace the previous red building. A description reads ”Pen+touch a different sketch-to-word to link it to the command. e) Subtitled ”(Confirmation — Tap again)” The diagram is shown with the updated yellow skyscraper sketch under the building, indicating that the re-linking was successful. The user taps the ”language action” button again to confirm. Below, the full scene in the canvas is shown with various building and skyscraper sketches and some clouds. A dashed-arrow between the ape and the final-selected skyscraper is shown to indicate that the command is about to make the ape jump on the skyscraper. A description reads, ”The ape jumps onto the final building.”
@@
[…]
[A screen-shot focused on the rectangular transcript view, its selection/de-selection mechanism, and associated buttons]A screen-shot focused on the rectangular transcript view, its selection/de-selection mechanism, and associated buttons. The user can toggle selection of ranges of words in the view with touch. The user’s finger is shown dragging from a word to another word to define a sub-range of words to remove from the command. The selected words are used to generate a command from the transcript view, so the de-selection of words will remove them. For the input sentence (that will be described now), ¡B¿ indicates the beginning position and ¡E¿ indicates the end position of the user’s finger, and — indicates new-line splits within the transcript view’s input: ”keep this part of the sentence ¡B¿ remove this part — of the sentence ¡E¿ keep this part of the sentence — remove this part of the sentence keep this part of the sentence”
@@
[…]
[Find by name functionality using the ”find panel”]Find by name functionality using the ”find panel”: a scene in the DrawTalking interface with many palm tree (”tree”) sketches, a lake sketch, and a sunset sketch. In the transcript view, the word ”tree” is selected, indicating a search for all objects named ”tree.” On the right, the find panel is shown containing miniature selectable entries corresponding to all the tree objects.
@@
[…]
[Toggling on/off states for user-defined rules using the rules search functionality in the ”find panel”]Toggling on/off states for user-defined rules using the rules search functionality in the ”find panel”: a scene containing a boy, dog, ball, and lake/pool/ The transcript view has the words, ”I can disable that too see so let’s look at the rules this is the water moving up one temporarily.” The word ”rule” in that sentence is highlighted, indicating that the find panel will show the rules. The panel has three entries for existing rules: 1: ”WHEN *ball collide with *water -¿ *water move for TIME THEN *water move for TIME”” is crossed-out, meaning the rule is inactive. 2: ”WHEN” *ball collide with *water -¿ I create *splash at *ball. 3: ”WHEN *splash appear -¿ I destroy *splash”
@@
4.4. Language Commands and Functionality
DrawTalking interprets the structure of language input into commands built from primitives. These primitives are intended to be a small sample of possible functionality that demonstrate our working concept. (They are partly inspired by existing software and design spaces (Shi et al., 2021; Dreams, 2020; Resnick et al., 2009).) (See the Info Sheet supplemental material.) We emphasize that our contribution is the interactive way of controlling elements. Our specific application and the fidelity and comprehensiveness of the supported behaviors are implementation details meant as a minimal demonstration of the DrawTalking concept. In other words, animation complexity and visual fidelity aren’t limitations inherent to the interface concept. These can be improved by implementers without rethinking the interaction design or contribution. We focus on DrawTalking’s expressiveness as a control mechanism for any potential behaviors.
Verbs are actions performable by sketches and the system, either built-in or user-defined in terms of other verbs by composition of existing primitives. Examples of implemented verb primitives include:
  • animations (e.g. move, follow, rotate, jump, flee)
  • state changes (e.g. create, transform)
  • events (e.g. collide with, press)
  • inequalities (e.g. equal, exceed)
@@ 4.4. Language Commands and Functionality
(like on, to, under) cause verbs to exhibit different behavior, e.g. ”The dog jumps on/under the bed” impacts the dog’s final position relative to the target. All verbs can use any such prepositions as input. Prepositions also might describe the spatial relationships between objects in a command (Figure 7).
@@
Nouns, pronouns, and deixis refer to object labels and are used to pick specific objects or specify types of objects. Using deixis while selecting an object will select the object immediately; pronouns can refer-back to objects, enabling commands that read less repetitively and more naturally: (e.g. ”The dog jumped. She jumped again.”)
The following choose 1+ specific objects for a command by matching object labels (nouns, optional adjectives, etc.) with words:
  • 1 or more objects: … ”the” ¡object label(s)¿ …
  • all objects: … ”all” ¡object label(s)¿ …
  • a number of objects: … ¡number¿ ¡object label(s)¿ …
  • a single random object: … ”a”/”an” ¡object label(s)¿ …
@@
Adjectives and adverbs are usable as labels that define the properties of objects. Verbs receive adjectives as arguments, which are evaluated continuously to modulate values and effects (e.g. magnitude, speed, size, distance, height). For example, ”fast”, ”slow,” ”excited” impact jump height and/or movement speed. Adjectives also disambiguate like-noun labeled-objects (e.g. ”first house” vs. ”second house”). Adjectives can be removed programmatically via negation, e.g. The thing is not fast.
Adverbs heighten adjective effects multiplicatively – e.g. ”very” and ”slightly” – and are chainable – e.g. ”very, very.” Special adjectives offer system-control: e.g. ”static” fixes sketches to the screen like a GUI element, useful for buttons, d-pads, score displays, etc.; ”visible”/”invisible” toggle sketch visibility by command (e.g. ”The ¡…¿ becomes visible/invisible”) or automatically when saying ”This is a(n) visible/invisible ¡thing…¿”.
@@
[…]
[Disambiguation of searched objects with same names, by scene hierarchy. The DrawTalking user interface is shown.]Disambiguation of searched objects with same names, by scene hierarchy. The DrawTalking user interface is shown. The canvas has a windmill base object, a windmill blade object (attached to the windmill base), a stone object, a blade (sword) object (attached to the stone). The user input in the transcript view (the command input) reads ”the blades on the windmill rotate;” The semantics diagram (command) shows: ”blade windmill rotate” Underneath the word ”blade” is a miniature of the blade object that is attached to the windmill. Underneath the word ”windmill” is a miniature of the windmill object. This all shows that the system correctly picked the blade object specified in the command for a blade on the windmill.
@@ 4.5. Examples
Figure 8. Windmill Simulation: Flexible process for constructing an interactive windmill built from user-defined rules, sketches, and triggers. The user can do this in any order and can try results at each step. This works on any sketch labeled ”blade.”
@@
[A process showing the construction of a windmill whose blades rotate as wind sketches spawn-in and move past the blades, all scripted-in using speech commands for defining rules, along with hand-drawn sketches]A process showing the construction of a windmill whose blades rotate as wind sketches spawn-in and move past the blades, all scripted-in using speech commands for defining rules, along with hand-drawn sketches; Panel 1 shows construction of the windmill with the base labeled ”windmill,” and the blades labeled ”blades.” The blades are attached to the windmill via the command (captioned below) ”I attach the blades to the windmill.” Panel 2 shows the user saving a hand-drawn wind spiral sketch so the system knows what to spawn when ”wind” is specified for a ”create” command. Panel 3 shows the definition of rules for collision between wind and blades, captioned: ”When wind collides with blades, blades rotate.” and ”After wind collides with blades, blades stop rotating.” Panel 4 shows that extra ”switch” and left-side wall sketches have been drawn. Pressing the switch causes wind to spawn from the left wall and move to the right. Several clones of the previously saved wind sketch are shown moving to the right and colliding with the windmill blades, causing them to spin clockwise. Captioned are the commands to define all of this: ”When I press the switch I create wind at the wall.” ”After wind appears, wind moves right.”
Figure 9. Pong into Breakout: Left: per-player paddles and points; Middle, Right: scene reoriented, second player’s objects removed, destructible bricks logic defined and objects added in-bulk into the desired region. (Note: not all commands shown.)
@@
[Steps for creation of a game of Pong that is then repurposed and extended using the same sketches into the similar game of Break-out]Steps for creation of a game of Pong that is then repurposed and extended using the same sketches into the similar game of Break-out: Panel 1: Pong contains 4 wall sketches and score counters for each of two players atop the walls. There is a start game switch in the bottom right that starts the game when pressed. Two paddles within the walls are being manipulated via touch controls. A ball sketch is shown bouncing towards the right paddle, which is about to reflect it left. The current score is 0-0. A version of the same scene is shown with object labels enabled and then a version with them turned off. The speech commands are written below: ”When balls collide with paddles, paddles reflect balls.”… ”When balls collide with first/second goals second/first scores increase.” (The slashes indicate these are two commands with differences in the words in the slashed positions.) Panel 2: additional steps transforming Pong into breakout. Objects corresponding to the second player have been removed so it becomes a one-player game. The scene is rotated vertically, a rectangular region is added on the top. Next the semantics diagram displays ”I pack region with *block” with the rectangle region in miniature under ”region.” Last the result of that command is shown where several blocks have been packed into the rectangular region and the region has been removed. (The point was to fill the game region with several breakable blocks.) The ball is shown being reflected off the paddle into the blocks, some of which have already been broken by this state. The score is 3. Panel 3: the final state with object names hidden. By now, the scene has been transformed into Breakout. The scene has been rotated vertically, and the second score counter has been deleted. The user has created several blocks filling the top part of the inner-walled area of the game area they’ve created. Some of the blocks have been destroyed by the user’s bouncing ball. Currently, the ball has bounced off the user’s paddle and is heading towards another breakable brick. The score counter reads 25 for the total number of points earned from breaking the blocks. All of these mechanics have been created from user-specified rules in speech commands.
@@
Figure 10. Day/Night Cycle with Transforming Villagers: A sun/moon transforms into the other in-time with collision events. When the moon appears, the villagers transform into ghosts, which chase the boy. This shows event-based behavior changes.
[…]
[Day/Night Cycle with Transforming Villagers - creation steps]Day/Night Cycle with Transforming Villagers - creation steps. Panel 1: sun, moon, cursed villager (labeled with the word ”curse” as a property), ghost sketches created. The user taps the save button, indicating all 4 sketches have been saved by name into the system for spawning later. Caption: ”Save sun, moon, villager, ghost; then delete moon and ghost.” Panel 2: There’s a left wall sketch (small rectangle), a cycle (circular shape) sketch, a right wall sketch (small rectangle. Arrows point from the left wall to the cycle to the right wall, and the right wall back to the cycle back to the left wall. These indicate a back-and-forth motion of the cycle between the left and right walls. The speech command is shown: ”When the cycle collides with the left wall, the sun transforms into a ”moon,” ”When the cycle collides with the right wall; Caption: the sun transforms into a ”sun.” (referring to the existing sun object.); Panel 3: top caption: ”distance between walls and speed of cycle determines rate of transformations — user can adjust these). The panel is divided left/right by a vertical line separator. On left is the sun, with three cursed villagers below. On right is the moon with three ghosts below, corresponding to the cursed villagers that will transform into ghosts when the moon appears. Arrows between left and right indicate the periodic changes between these two scene states due to the aforementioned logic. Below read the two commands, ”When moons appear, all cursed villagers transform in ghosts” and ”When suns appear, all ghosts transform into cursed villagers.” Panel 4. Three ghost sketches chase a boy sketch, the moon is shown, indicating it’s the night-time state. A right arrow pointing from the boy indicates that the user might move the boy around. Arrows from the ghosts towards the boy indicate their following-the-boy behavior. The speech command for completing this set-up reads, ”When ghosts appear, ghosts follow the boy”
@@
Figure 11. ”Dog Plays Fetch” as a Prototype Game Mechanic (demonstrating adjectives for dynamic logic): The dog interaction is re-imagined as an interactive game mechanic.
[…]
[Constructing a game mechanic in which a button causes a dog to collect point-giving treat objects near a main character sketch.]Constructing a game mechanic in which a button causes a dog to collect point-giving treat objects near a main character sketch. The scene shows a circular sketch named ”proximity” moving to the center of a ”boy” sketch, corresponding to the command described in description a): ”The proximity teleports to the boy and then it attaches to him and disappears” (Auto-alignment and attachment of a custom invisible collider; in the scene are several objects labeled ”treat”, a button sketch, and a dog sketch. The treats colliding with the circular proximity collider described before receive the label ”nearby” as a result of the rules defined in description b): ”When the proximity collides with treats, treats become nearby.”, ”After the proximity collides with treats, treats become not nearby” (Change properties of collided objects). Description c) for another command for this setup reads: ”When I press the button, the dog moves to nearby treats.” (In this gameplay, the dog is constrained to move only to treats that are labeled ”nearby” by the boy’s proximity collider.) An arrow points from the dog to a treat labeled ”nearby” within the boy’s proximity collider radius, with the user’s finger tapping the aforementioned button, indicating that the dog has been commanded by the rule to move to a treat that is nearby. There is also a score (number) sketch labeled ”score,” mentioned in description d) for the rule: ”When the dog collides with treats, the score increases and then the dog destroys treats.” (Collectibles with working points mechanic with number sketch.) The score sketch is shown increasing from 0.00 to 1.00 by the dog collecting a nearby treat.
@@ 4.5.5. Day/Night Cycle: Transforming Objects and Periodic Collision Logic
5. Implementation
5.1. Hardware and Software
We implemented DrawTalking natively on the iPad Pro M2, with the main code written in C++ and C, bridged with Objective C++ and Swift to access platform-specific APIs. Speech recognition is on-device. A local NodeJS+Python server-side for NLP runs on a MacBook Pro. Text is continuously sent to the server where 1) a dependency-tree is created, 2) sent to the client and compiled into a key-value format for semantic roles (e.g. AGENT, OBJECT), 3) either interpreted in real-time by our custom engine to drive interactive simulations, or used to find objects by deixis. In final stages, the system looks-up and maps objects to roles in the command structure and executes it to produce animations, effects, and state changes. (See Figure 12 for a diagram of the the concrete client-server implementation and data-flow.)
5.2. Processing Steps
The language processing component comprises roughly 4 steps that translate a natural language data structure into a command for the application (Figure 13). These user-initiated commands are scripts constructed and executed at run-time. As they’re executed, the scripts handle data states, sequences, and dynamic instantiations of scripts corresponding to verbs in our built-in library. Verb scripts are modules comprising arbitrary code and accepted parameters, as opposed to generated code. The system draws from ideas in visual programming and virtual machine-based engines (cycling74, 2018; Resnick et al., 2009; Another World, [n. d.]; Little Big Planet 2, 2011). Note that implementers can add functionality without modifying the rest of the system. For supplementary examples, see Appendix D.